OpenStudy (anonymous):

fan and medal given to best answer !! Find the difference of (−5x^2 − 4x^3 − 3) − (9x^3 + 3x − 8x^2). Be sure to show all of your work in a neat and organized fashion.

OpenStudy (nurali):

−5x^2 − 4x^3 − 3) − (9x^3 + 3x − 8x^2) = −5x^2 − 4x^3 − 3 − 9x^3 − 3x + 8x^2 Combine like terms: = (−4 − 9)x^3 + (−5 + 8)x^2 − 3x − 3 = −13x^3 + 3x^2 − 3x − 3

OpenStudy (anonymous):

(−5x2 − 4x3 − 3) − (9x3 + 3x − 8x2) = (−5x2 − 4x3 − 3) + −1(9x3 + 3x − 8x2)

OpenStudy (anonymous):

distribute the -1 then remove the brackets and find the difference

OpenStudy (anonymous):

combine like terms